# Diffusion Dataset Condensation: Training Your Diffusion Model Faster with Less Data

### v2

@@ 3.2 Attach: Semantic and Visual Information Enhancement
To further enhance the information richness of the selected data, we introduce the Attach phase that augments each instance with semantic and visual information. While the Select phase generates a compact and informative subset, the achievable performance ceiling based solely on this phase remains limited. Consequently, we inject more comprehensive semantic and visual representations into the selected subset to further bolster our method’s generalization capability.
Dual Conditional Embedding (DC-Embedding). Existing C2I synthesis methods [11, 13] commonly rely on class embeddings trained from scratch, which often fail to effectively capture inherent semantic information (see Appendix F.1). We enrich the class embedding by incorporating text representations derived from a pre-trained text encoder (e.g., T5-encoder [35]). For each class $c\in\{1,\dots,C\}$, a descriptive prompt $P(c)$ (e.g., “a photo of a cat”) is encoded by a pre-trained text encoder $f_{\text{text}}$, yielding its corresponding text embedding $t_{c}$ and text mask $t_{\textrm{mask}}$:
$\begin{split}t_{c},t_{\textrm{mask}}&=f_{\text{text}}(P(c)),\\ \end{split}$ (8)
@@
[…]
Figure 4: Visualization of samples from Random (i.e., random sampling) vs. D2C. D2C demonstrates a significant performance advantage over Random across all training iterations, under both a strict 0.8% (10K) and a 4% (50K) data budget, leading to a substantial acceleration in training.
Figure 5: Overview of DC-Embedding.
@@
[…]
@@ 3.3 D2C Training Process
Here, we detail the training process of the diffusion model using our condensed dataset, which comprises a compact subset selected during the Select phase and subsequently enriched with semantic and visual information during the Attach phase. Our goal is to fully leverage the information existed within our condensed dataset to accelerate training without compromising performance.
We employ a conditional diffusion model $\mathcal{D}_{\theta}$ and, as an example, utilize the optimization objective of score-based diffusion models: predicting the added noise $\epsilon$ from the perturbed latent input $\mathbf{x}_{t}$ at time step $t$, conditioned on the text information $y_{\textrm{text}}$ and the class label $y$. The new denoising loss is defined as $\mathcal{L}^{\text{new}}_{\text{diff}}=\mathbb{E}_{\mathbf{x}_{0}\sim q_{0}(\mathbf{x}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t,y,y_{\textrm{text}})\|^{2}_{2}\right]$, where the specific injected forms of $y$ and $y_{\text{text}}$ can be found in Sec. 3.2. Then, to maximize the utilization of visual information, we adopt the same formulation as REPA [12], which involves aligning the encoder’s output (i.e., the decoder’s input) within the diffusion model with the visual representation $y_{\text{vis}}=\{v_{i}\}_{i=1}^{h}$. Concretely, from a designated intermediate layer of the diffusion backbone, we obtain token features $\{h_{i}\in\mathbb{R}^{d}\}_{i=1}^{h}$. A projection head $\phi$ maps these tokens from $\mathbb{R}^{d}$ to $\mathbb{R}^{d_{\text{text}}}$, and we compute a semantic alignment loss:
@@
This loss encourages the model to align its encoder’s output with visual representations extracted from the visual encoder, promoting localized realism and spatial consistency [36] in generation.
@@
This unified training strategy enables D2C to effectively learn from limited yet enhanced data, offering a practical solution for efficient diffusion training under resource-constrained settings.
Table 1: Comparison of FID-50K across various dataset condensation methods and data budgets using DiT-L/2 and SiT-L/2 on ImageNet 256$\times$256. We use CFG=1.5 for evaluation. D2C surpasses other methods at all settings.
@@
Table 2: Comparison with a strict data budget 0.8% (10K) on ImageNet 512$\times$512. We use CFG=1.5 for evaluation. D2C surpasses random sampling at all settings. D2C’s performance on SiT-L/2 with 300k iterations can be found in Appendix J.
@@
DiT-L/2 Random 100k 24.8 11.9 74.3 0.65 0.42
DiT-L/2 D2C (Ours) 100k 14.8 6.9 109.2 0.63 0.52
DiT-L/2 Random 300k 17.1 12.8 130.6 0.64 0.41
DiT-L/2 D2C (Ours) 300k 5.8 15.1 318.9 0.77 0.29
SiT-L/2 Random 100k 13.3 22.8 197.1 0.69 0.68
SiT-L/2 D2C (Ours) 100k 9.1 14.3 261.7 0.72 0.34
4 Experiments
We validate the performance of D2C and analyze the contributions of its components through extensive experiments. In particular, we aim to answer the following questions:
1) Can D2C improve training speed and reduce data usage of diffusion models? (Table 1, Figs. 14)
2) Does D2C generalize well across backbones, data scales, and resolutions? (Tables 12, Fig. 4)
3) How do D2C’s components and hyperparameter choices affect its overall effectiveness? (Figs. 368)
Experiment settings. We conduct experiments on the ImageNet-1K dataset [37], using subsets of 10K, 50K, and 100K images, corresponding to 0.8%, 4%, and 8% of the full dataset, respectively. All images are center-cropped and resized to 256$\times$256 and 512$\times$512 resolutions using the ADM [16] preprocessing pipeline. Furthermore, we use [$\cdot$]-L/2 and [$\cdot$]-XL/2 architectures in both DiT [11] and SiT [13] backbones, following the standard settings outlined in Ma et al. [13]. More details on implementation and training can be found in Appendix C.
Figure 6: Left: Ablation studies of the interval value $k$ in interval sampling at different data budgets. With a small interval value, training can experience significant acceleration during the early iterations. However, due to the limited diversity of the selected dataset, performance tends to be overtaken by random sampling in later iterations. As the interval value increases, the final FID-10K scores initially improve and then decline, suggesting an optimal interval value exists. Notably, this optimal value is 96 for a 10K data budget and 16 for a 50K data budget, and the ratio 96/16 closely approximates the ratio of the data budgets, 50K/10K. Right:. Ablation study of DC-Embedding under a 10K data budget. “Only Class” stands for the baseline approach of injecting class embeddings. This figure demonstrates that both the text embedding branch and the original class embedding branch (see Fig. 5) are critical for the effective training of diffusion models.
Table 3: D2C vs. SRe2[24] on ImageNet 256$\times$256 with a strict data budget 0.8% (10K).
@@ 4 Experiments
DiT-L/2 SRe2L 104.2 20.2 14.1 0.20 0.30
DiT-L/2 D2C (Ours) 4.2 11.0 283.6 0.72 0.24
SiT-L/2 SRe2L 82.3 19.8 18.1 0.27 0.19
SiT-L/2 D2C (Ours) 3.9 10.7 289.7 0.73 0.25
Evaluation and baselines. We train models from scratch on the collected subset and evaluate them using FID [38], sFID, Inception Score [39], Precision (Prec.), and Recall (Rec.), adhering to standard evaluation protocols [16, 11, 13]. We compare our method against REPA [12], and various data condensation and selection baselines, including SRe2[24], Herding, K-Center, and random sampling, using SiT and DiT architectures [13, 11]. Further details regarding evaluation metrics and baseline methods can be found in Appendix G and H.
4.1 Main Result
Training Performance and Speed. We evaluate D2C on SiT-XL/2 using 10K and 50K data budgets, comparing its performance against REPA and a vanilla SiT model trained on the full ImageNet dataset (a 1.28M data budget), as well as random selection with 10K and 50K data budgets. As illustrated in Fig. 1 (b), our method achieves a FID-50K of 4.23 at only 40K iterations with 10K training data and without CFG. In contrast, REPA requires 4 million steps and the vanilla SiT model needs over 7 million steps to reach comparable performance, representing an acceleration of over 100$\times$ and 233$\times$, respectively. Under a 4% data budget (50K) with CFG set to 1.5, our method achieves an FID of 2.78 at 180K steps, further demonstrating significant data and compute efficiency (Fig. 1 (c)). Moreover, Fig. 4 presents a visual comparison between random selection and our D2C at 10K and 50K data sizes, where images are generated from the same noise and class label. Our method demonstrates superior visual quality compared to the baseline and generates higher-quality images, even during the early iterations of training.
Comparison on ImageNet 256$\times$256. We compare D2C with random sampling, Herding [40], K-Center [30], and SRe2[24] under various data budgets and backbones. As illustrated in Table 1, D2C consistently achieves the lowest FID across all settings. For instance, using only 0.8% of the data and 100K iterations with early stopping, our method achieves a FID-50K of 4.20 on DiT-L/2 and 3.98 on SiT. These results demonstrate the superiority of our approach over existing methods. Notably, SRe2L, which performs well in classification task, fails on this generative task (see Table 3) due to its focus on category-discriminative features. Similarly, geometry-based methods like Herding and K-Center, along with random sampling, prove inadequate for achieving efficient and high-performing training of diffusion models.
Comparison on ImageNet 512$\times$512. As shown in Table 2, D2C achieves a FID of 5.8 on DiT-L/2, a significant improvement over the 17.1 achieved by random sampling at 300k iterations under the ImageNet 512$\times$512 settings. On SiT-L/2, similar improvements are observed. These results demonstrate that D2C generalizes well to higher resolutions.
@@ 4.2 Ablation Study
Figure 7: Ablation sutdies of interval sampling and DC-Embedding. “Mean=N” denotes the selection of a subset of samples where the average diffusion loss is closest to the value N.
Ablation on Select Phase. We investigate the impact of the interval value $k$ in the Select phase, as shown in Fig. 6 (Left) and Fig. 7. Using a small value accelerates early training by prioritizing min-loss samples, which are simpler and easier to learn. However, the limited diversity of such samples leads to degraded performance in later stages, eventually being overtaken by settings with moderate interval values. In contrast, large intervals or random selection introduce excessive max-loss or uncurated samples, destabilizing training (Fig. 3). As $k$ increases, we observe that FID-10K first decreases and then worsens, revealing an optimal trade-off between diversity and learnability. Empirically, the best results are achieved with an interval of 96 for the 10K budget and 16 for 50K, approximately following the ratio of data budgets (50K/10K).
@@
Table 4: Ablation on semantic and visual information in the Attach phase.
Ablation on Attach Phase. We conduct ablation studies on the Attach phase from two perspectives. First, as shown in Fig. 6 (Right), we examine the effectiveness of dual conditional embedding under a 10K data budget. Using only class or text embedding results in degraded performance compared to their combination, indicating the necessity of both components to capture category semantics and text embeddings. Notably, text embedding alone performs better than class embedding, suggesting that textual descriptions provide richer semantic information for generative tasks. Second, as shown in Table 4, we analyze the impact of the semantic and visual information injection modules. The FID-10K increases to 37.07 when both branches are removed and to 9.01 when only the visual injection module is excluded. Notably, the combination of both modules yields the best FID of 7.62. These results underscore the crucial role of injecting both semantic and visual information in accelerating diffusion model training.
@@ 5 Conclusion
In this paper, we introduce D2C, the first dataset condensation franework that can significantly accelerate diffusion model training for the generative task. Our pipeline comprises two key phases: Select and Attach. In the Select phase, we leverage diffusion difficulty score and interval sampling to obtain a subset that is both compact and diverse. Subsequently, in the Attach phase, we augment this subset with critical semantic and visual information, leading to a dramatic improvement in the training acceleration ratio and enhanced performance robustness. While there are still some limitations (refer to Appendix A), our D2C as the pioneering work in this novel approach still achieved impressively 100233×100$\sim$ 233$\times$100 ∼ 233 × training acceleration over the baselines. We believe that our work will offer novel inspirations and motivate more research on this promising field in the future.
Broader Impact
We introduce D2C, a method engineered to decrease training overhead while minimizing performance degradation. The dataset employed for training D2C was ImageNet-1K, which adheres to ethical guidelines. Moreover, we prioritize the responsible and ethical deployment of this technology, aiming to maximize its societal benefits while proactively addressing and mitigating any potential risks.
Appendix
Appendix A Limitation
This work primarily focuses on accelerating and improving the training process of diffusion models. While our approach demonstrates promising results in the early stage, the final performance upon convergence does not yet fully match baseline benchmarks. It is important to unleash the potential of our algorithm, such as fine-tuning the vision and text encoders to better align noise latent features with the integrated visual information, thereby reducing existing discrepancies. Additionally, although the current study centers on image generation tasks, extending our method to other domains, such as 3D generation and video synthesis, represents a promising direction for future exploration.

### v3

@@ 3.2 Attach: Semantic and Visual Information Enhancement
To complement the Select phase, which yields a compact subset of informative real images, the Attach phase enriches each selected instance with additional semantic and visual information. In particular, we attach semantic information via a Dual Conditional Embedding (DC-Embedding) module and inject visual information through visual representation, resulting in a more expressive condensed dataset and improved generalization of the trained diffusion models.
Dual Conditional Embedding (DC-Embedding). Existing C2I synthesis methods [33, 28] commonly rely on class embeddings trained from scratch, which often fail to effectively capture inherent semantic information (see Appendix I.1). We enrich the class embedding by incorporating text representations derived from a pre-trained text encoder (e.g., T5-encoder [31]). For each class $c\in\{1,\dots,C\}$, a descriptive prompt $P(c)$ (e.g., “a photo of a cat”) is encoded by a pre-trained text encoder $f_{\text{text}}$, yielding its corresponding text embedding $t_{c}$ and text mask $t_{\textrm{mask}}$:
[…]
@@
$\tilde{t}_{c}$ $=\text{Conv1d}(t_{c}\times t_{\textrm{mask}}),\quad y_{\text{text}}=\text{MLP}(\tilde{t}_{c})+\tilde{t}_{c}+e_{c}.$ (9)
Figure 5: D2C improves visual quality under tight data budgets. We compare Random sampling and D2C on DiT-L/2 at 10k and 50k data budgets, and neither setting uses classifier-free guidance.
@@
where $N$ is the number of image patches and $d_{\text{text}}$ is the feature dimension. We retain the first $h$ (i.e., number of tokens in the diffusion transformer) tokens of $y_{\text{vis}}$ to form a compact representation of the dominant structure: $y_{\text{vis}}=y_{\text{vis}}[:\!h,:]\in\mathbb{R}^{h\times d_{\text{text}}}$. As outlined in REPA [53], this visual information provides a semantic prior for the diffusion model and thus significantly benefits data-centric efficient training. Similar to the text information $y_{\text{text}}$, the visual information $y_{\text{vis}}$ is also stored on disk as attached metadata alongside the selected subset $\mathcal{D}_{\text{IS}}$.
@@ 3.3 D2C Training Process
Here, we detail the training process of the diffusion model using our condensed dataset, which comprises a compact subset selected during the Select phase and subsequently enriched with semantic and visual information during the Attach phase. Our goal is to fully leverage the information contained in our condensed dataset to accelerate training without compromising performance.
We employ a conditional diffusion model $\mathcal{D}_{\theta}$ and, as an example, utilize the optimization objective of score-based diffusion models: predicting the added noise $\epsilon$ from the perturbed latent input $\mathbf{x}_{t}$ at time step $t$, conditioned on the text information $y_{\textrm{text}}$ and the class label $y$. The new denoising loss is defined as $\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{x}_{0}\sim q_{0}(\mathbf{x}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t,y,y_{\textrm{text}})\|^{2}_{2}\right]$, where the specific injected forms of $y$ and $y_{\text{text}}$ can be found in Sec. 3.2. Then, to maximize the utilization of visual information, we adopt the same formulation as REPA [53], which involves aligning the encoder’s output (i.e., the decoder’s input) within the diffusion model with the visual representation $y_{\text{vis}}=\{v_{i}\}_{i=1}^{h}$. Concretely, from a designated intermediate layer of the diffusion backbone, we obtain token features $\{h_{i}\in\mathbb{R}^{d}\}_{i=1}^{h}$. A projection head $\phi$ maps these tokens from $\mathbb{R}^{d}$ to $\mathbb{R}^{d_{\text{text}}}$, and we compute a semantic alignment loss:
@@
This loss encourages the model to align its encoder’s output with visual representations, promoting localized realism and spatial consistency [32] in generation.
@@
This training strategy enables D2C to effectively learn from limited yet enhanced data, offering a practical solution for efficient diffusion training under minimal budgets.
4 Experiments
In this section, we validate the performance of D2C and analyze the contributions of its components through extensive experiments. In particular, we aim to answer the following questions: 1) Can D2C improve training speed and reduce data usage of diffusion models? 2) Does D2C generalize well across backbones, data scales, and resolutions? 3) How do D2C’s components and hyperparameter choices affect its overall effectiveness?
Table 1: Comparison of gFID-50K across various dataset condensation methods and data budgets using DiT-L/2 and SiT-L/2 on ImageNet $\times256$. We use CFG=1.5 for evaluation. D2C surpasses other methods at all settings.
@@ 4 Experiments
4.1 Setup
Experiment settings. We conduct experiments on the ImageNet-1K dataset [36], using subsets of 10K, 50K, and 100K images, corresponding to 0.8%, 4%, and 8% of the full dataset, respectively. To further demonstrate the generalization and effectiveness of our method, Appendix J reports additional results of D2C on CIFAR datasets. All images are center-cropped and resized to $\times256$ and $\times512$ resolutions using the ADM [6] preprocessing pipeline. Furthermore, we use $\cdot$]-L/2 and $\cdot$]-XL/2 architectures in both DiT [33] and SiT [28] backbones, following the standard settings outlined in Ma et al. [28].
Table 2: Comparison with a strict data budget 0.8% (10K) on ImageNet $\times512$. We use CFG=1.5 for evaluation. D2C surpasses random sampling at all settings.
@@ 4.1 Setup
DiT-L/2 Random 100k 24.8 11.9 74.3 0.65
DiT-L/2 D2C (Ours) 100k 14.8 6.9 109.2 0.63
DiT-L/2 Random 300k 17.1 12.8 130.6 0.64
DiT-L/2 D2C (Ours) 300k 5.8 15.1 318.9 0.77
SiT-L/2 Random 100k 13.3 22.8 197.1 0.69
SiT-L/2 D2C (Ours) 100k 9.1 14.3 261.7 0.72
SiT-L/2 Random 300k 5.0 13.6 316.9 0.76
SiT-L/2 D2C (Ours) 300k 4.22 11.6 289.7 0.79
Evaluation and baselines. We train models from scratch on the collected subset and evaluate them using gFID [16], sFID, Inception Score [37] and Precision, adhering to standard evaluation protocols [6, 33, 28]. We compare our method against REPA [53], REPA-E [22], REG [48] and various data condensation and selection baselines, including SRe2[52], RDED [44], Herding, K-Center, and random sampling, using SiT and DiT architectures [28, 33]. Further details regarding evaluation metrics and baseline methods can be found in Appendix D and E.
Figure 6: Left: Interval-sampling ablation. Small $k$ speeds early training. The best final gFID-10K appears at k=96k{=}96 for the 10K budget and k=16k{=}16 for the 50K budget, roughly scaling with data size. Right: DC-Embedding ablation at 10K. Combining text and class embeddings outperforms either alone; “Only Class” denotes the baseline that injects class embeddings only.
4.2 Main Result
@@
Table 3: Comparison of acceleration algorithms on ImageNet-1K.
[…]
DiT L-2 1.28M 400k 23.3
++ REPA 1.28M 400k 15.6
++ D2C 0.05M 10k 14.81
++ D2C 0.01M 10k 4.2
SiT L-2 1.28M 400k 18.8
++ REPA 1.28M 700k 8.4
++ D2C 0.01M 80k 7.07
SiT XL-2 1.28M 7M 8.3
++ REPA 1.28M 4M 5.9
++ REPA-E 1.28M 235k 5.9
++ REG 1.28M 200k 5.0
++ D2C 0.01M 40k 4.3
++ D2C 0.05M 180k 2.78
Comparison on ImageNet $\times256$. We compare D2C with random sampling, Herding [5], K-Center [18], SRe2[52], RDED [44] under various data budgets and backbones. As shown in Table 1, D2C consistently achieves the lowest FID across all settings. For instance, using only 0.8% of the data and 100K iterations with early stopping, our method achieves a gFID-50K of 4.20 on DiT-L/2 and 3.98 on SiT. These results demonstrate the superiority of our approach over existing methods. Notably, SRe2L and RDED, which perform well in classification task, fail on this generative task (see Table 4) due to their focus on category-discriminative features. Similarly, geometry-based methods like Herding and K-Center, along with random sampling, prove inadequate for achieving efficient and high-performing training.
Comparison on ImageNet $\times512$. As shown in Table 2, D2C achieves a gFID of 5.8 on DiT-L/2, a significant improvement over the 17.1 achieved by random sampling at 300k iterations under the ImageNet $\times512$ settings. On SiT-L/2, similar improvements are observed. These demonstrate that D2C generalizes well to higher resolutions.
@@ 4.3 Ablation Study
Table 4: D2C vs. SRe2[52] and RDED [44] on ImageNet $\times256$ with a data budget 0.8% (10K).
[…]
DiT-L/2 RDED 166.2 60.1 10.8 0.09
DiT-L/2 SRe2L 104.2 20.2 14.1 0.20
DiT-L/2 D2C (Ours) 4.2 11.0 283.6 0.72
SiT-L/2 RDED 97.5 66.8 65.63 0.22
SiT-L/2 SRe2L 82.3 19.8 18.1 0.27
SiT-L/2 D2C (Ours) 3.9 10.7 289.7 0.73
Ablation on Select Phase. We investigate the impact of the interval value $k$ in the Select phase, as shown in Fig. 6 (Left). Using a small value accelerates early training by prioritizing min-loss samples, which are simpler and easier to learn. However, the limited diversity of such samples leads to degraded performance in later stages, eventually being overtaken by settings with moderate interval values. In contrast, large intervals or random selection introduce excessive max-loss or uncurated samples, destabilizing training (Fig. 3). As $k$ increases, we observe that gFID-10K first decreases and then worsens, revealing an optimal trade-off between diversity and learnability. Empirically, the best results are achieved with an interval of 96 for the 10K budget and 16 for 50K, approximately following the ratio of data budgets (50K/10K). Table 5 further shows that using the Select stage alone reduces gFID from 37.07 to 14.96, underscoring its effectiveness and usefulness.
Ablation on Attach Phase. We evaluate Attach from two angles. First, as shown in Fig. 6 (Right), DC embedding consistently outperforms using either alone under a 10K budget, with text-only better than class-only, indicating richer semantics from textual descriptions. Second, Table 5 shows steady gains from the injection modules: baseline gFID-10K is 14.96, adding only visual information reaches 10.37, adding only DC embedding reaches 9.01, and combining both achieves the best 7.62. Appendix I.2 further ablates the visual encoder and demonstrates the robustness of our approach.
Table 5: Ablation studies on the Select and Attach phases. Sel.: Select. Vis.: Vision.
@@
Table 6: A breakdown of the computational overhead for sub-processes in D2C. Compared to the REPA baseline, the additional scoring time is negligible, demonstrating D2C’s efficiency.
[…]
REPA N/A N/A 4M 750h 5.9
D2C (w/o select) N/A N/A 0.04M 7.4h 5.6
D2C (w/ select) From Scratch 1.9h 0.04M (7.4+26.2)h 4.9
D2C (w/ select) Pretrained 2.1h 0.04M 7.4h 4.3
[…]
@@ 5 Conclusion
In this paper, we introduce D2C, the first dataset condensation framework that significantly accelerates diffusion model training for generative tasks. D2C follows a two-phase pipeline, Select and Attach, which selects a compact yet diverse subset via a diffusion difficulty score with interval sampling and enriches it with semantic and visual signals. On ImageNet-1K, D2C achieves $100\text{--}233\times$ faster training than strong baselines while maintaining competitive generative quality, and we hope it will motivate further research on data-centric efficiency for diffusion models.
Acknowledgement. This work was supported by the National Natural Science Foundation of China under Grant No. 62506317.
Supplementary Material
Appendix A Positioning D2C within Dataset Condensation Paradigms
While some works equate dataset condensation with gradient-based pixel-level optimization of synthetic images, a broader line of literature defines it as constructing compact training sets that retain the learning efficacy of the original data [49], which also includes image-level schemes such as OD3[20] and RDED[44]. In this broader paradigm, the key objective is not how the condensed data are obtained, but whether the resulting small dataset can support training models that closely match the performance of those trained on the full dataset. D2C follows this latter view. It condenses the dataset by selecting a highly informative subset guided by diffusion difficulty and then attaching additional semantic and visual information that enriches each sample without altering its raw pixels. This design is analogous in spirit to OD3 and RDED, which also operate at the level of image selection rather than direct pixel optimization. Consequently, D2C naturally fits within the dataset condensation family, while being specifically tailored to generative diffusion models and addressing a gap that is not covered by existing pixel-level condensation methods.